A train leaves a city heading west and travels at 50 mph. Three hours later, a second train leaves from the same place and travels in the same direction at 65 mph. How long will it take for the second train to catch up with the first train?
Answer by ewatrrr(24785)


Hi,
Let t represent the total time the train traveling 50mph takes before the
second train catches up. (t-3) would represent the time for the second train
Distances for each train are the same.
r*t=d ****
50mph*t = 65mph(t-3)
50mph*t = 65mph*t - 3*65mph
3*65mph = 15mph*t
3*65mph/15mph = 13hrs, the time the first train travels
Second train travels for 10hrs (13hr-3hr)
checking our answer
50*13 = 650= 65*10